9) Use data from Appendix IIB in the textbook to calculate the equilibrium constants at 25∘C for each of the following reactions.
part a) 2NO(g)+O2(g)⇌2NO2(g)
part b) N2(g)+O2(g)⇌2NO(g)

10) Consider the following reaction:
CO(g)+2H2(g)⇌CH3OH(g)
Kp=2.26×104 at 25 ∘C. Calculate ΔGrxn for the reaction at 25 ∘C under each of the following conditions.
part a) standard conditions
part b) at equilibrium
part c) PCH3OH= 1.5 atm ;
PCO=PH2= 1.2×10−2 atm
1
Expert's answer
2020-04-18T06:45:29-0400

(10)We know "\u2206G_{rxn}=-RTln(K)"

(a)at standard conditions

"\u2206G_{rxn}=-8.314\u00d7298\u00d7ln(2.26\u00d710^4)"

"\u2206G_{rxn}=-8.314\u00d7298\u00d71.8345=4545.13232"

(b)at equilibrium

"\u2206G=0"

(c)"P_{CH_3OH}=1.5 atm;P_{CO}=0.012atm;P_{H_2}=0.012atm"

"K_p=\frac{[CH_3OH]}{[CO][H_2]^2}=\frac{1.5}{0.012^3}=8.68×10^{5}"

"\u2206G_{rxn}=-8.314\u00d7298\u00d7ln(8.68\u00d710^5)"

"\u2206G_{rxn}=-8.314\u00d7298\u00d75.94"

"\u2206G_{rxn}=14713.1102"

(9)(a)"2NO+O_2->2NO_2"

0.002 0.001 0.05

At equilibrium no of moles of each compound.

"K_c=\frac{[NO_2]^2}{[NO]^2[O_2]}"

"K_c=\frac{[0.05]^2}{[0.002]^2[0.001]}=6.9×10^5/mol"

(b)"N_2(g)+O_2(g)->2NO(g)"

0.036 0.0089 3.6

At equilibrium no of moles of each compound.

"K_c=\frac{[NO]^2}{[N_2][O_2]}"

"K_c=\frac{[3.6]^2}{[0.036][0.0089]}=4.1×10^4"